### v1

@@ 5 Conclusion
In this work, we have investigated the stochastic 3D Navier-Stokes equation in the presence of thermal noise, which corresponds to model A of FNS, using both direct numerical simulations and functional renormalisation group. We have identified a new fixed point, called IFNS, found in the inviscid (Euler) limit of model A, and which controls the large momentum behaviour of the correlation function. This fixed point is characterised by a critical dynamical exponent $z=1$ at small time delays. Moreover, the FRG result shows that the behaviour of the temporal correlations crossover at large time delays to a much slower decay with $z=2$. This second time regime had remained elusive in direct numerical simulations so far due to the difficulty to properly resolve long time delays. Our numerical simulations of model A have allowed us to observe both regimes, yielding the first numerical evidence of the long-time regime for 3D Navier-Stokes equation. The simulations have thus fully confirmed the FRG predictions.
@@
Remarkably, the unveiled scaling $z=1$ is common to the inviscid regime of both the Burgers and Navier-Stokes equations. Moreover, the value of $z$ is independent of the dimension, as well as of the precise correlations of the stochastic term $\mathbf{f}$ provided it has a Gaussian distribution [34, 55]. Thus it holds both for equilibrium situations (when $\mathbf{f}$ is a thermal noise) and non-equilibrium situations (when $\mathbf{f}$ is a large-scale forcing), even though the corresponding static behaviours, eg. the kinetic energy spectra, are completely different. We have provided the explanation of this robustness, as we have shown that the large momentum behaviour of correlation functions is completely fixed and solely determined by the symmetries. Indeed, all these hydrodynamical equations share the same fundamental (extended) symmetries: the time-dependent Galilean symmetry and time-dependent shifts of the response fields (although the latter bears slightly different forms in Burgers and NS equations). These symmetries, through the related Ward identities, allow for an exact closure of the FRG flow equations in the asymptotic regime of large momenta. Their solution endows identical forms for Burgers or NS flows, differing only by non-universal constants, and featuring the exact $z=1$ exponent. The FRG has thus proved very efficient to understand the temporal behaviour of correlation functions. In perspective, it would be valuable to use it to gain more understanding of static properties of structure functions, and in particular intermittency effects.
Acknowledgements
@@
LC wishes to thank G. Eyink for suggesting this work. LG acknowledges support by the MSCA Cofund QuanG (Grant Number : 101081458) funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or Université Grenoble Alpes. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 6 Conclusion
In this work, we have investigated the stochastic 3D Navier-Stokes equation in the presence of thermal noise, which corresponds to model A of FNS, using both direct numerical simulations and functional renormalisation group. The aim was to explore what happens when taking the inviscid limit in this model, which had been studied so far only in the perturbative regime, corresponding to large viscosities. We have indeed identified a new non-perturbative, strong-coupling, fixed point, called IFNS, found in the inviscid (Euler) limit of model A, and which also controls, when the viscosity is finite, the large momentum behaviour of the correlation function. This fixed point is characterised by a critical dynamical exponent $z=1$ at small time delays. Moreover, the FRG result shows that the behaviour of the temporal correlations crosses over at large time delays to a much slower decay with $z=2$. This second time regime had remained elusive in direct numerical simulations so far due to the difficulty to properly resolve long time delays. Our numerical simulations of model A have allowed us to observe both regimes, yielding the first numerical evidence of the long-time regime for 3D Navier-Stokes equation. The simulations have thus fully confirmed the FRG predictions. This work has brought a further theoretical insight in the behaviour of the stochastic Navier-Stokes equation, although the IFNS scaling may not be observable in a real fluid at equilibrium.
@@
Remarkably, the unveiled $z=1$ scaling is common to the inviscid regime of both the Burgers and Navier-Stokes equations. Moreover, the value of $z$ is independent of the dimension, as well as of the precise correlations of the stochastic term $\mathbf{f}$ provided it has a Gaussian distribution [38, 60]. Thus it holds both for equilibrium situations (when $\mathbf{f}$ is a thermal noise, and the stationary state is a fluid at rest) and non-equilibrium situations (when $\mathbf{f}$ is a large-scale forcing, and the stationary state is a turbulent fluid), even though the corresponding static behaviours, eg. the kinetic energy spectra, are completely different. We have provided the explanation of this robustness, as we have shown that the large momentum behaviour of correlation functions is completely fixed and solely determined by the symmetries. Indeed, all these hydrodynamical equations share the same fundamental (extended) symmetries: the time-dependent Galilean symmetry and time-dependent shifts of the response fields (although the latter bears slightly different forms in Burgers and NS equations). These symmetries, through the related Ward identities, allow for an exact closure of the FRG flow equations in the asymptotic regime of large momenta. Their solution endows identical forms for Burgers or NS flows, differing only by non-universal constants, and featuring the exact $z=1$ exponent. The FRG has thus proved very efficient to understand the temporal behaviour of correlation functions. In perspective, it would be valuable to use it to gain more understanding of static properties, encoded for instance in structure functions, and in particular of intermittency effects.
[…]
@@
LC wishes to thank G. Eyink for suggesting this work and for his valuable comments about the physical relevance of the IFNS scaling. LG acknowledges support by the MSCA Cofund QuanG (Grant Number : 101081458) funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or Université Grenoble Alpes. Neither the European Union nor the granting authority can be held responsible for them.